# Banach Poisson–Lie groups, Lax equations and the AKS theorem in infinite dimensions

### v1

@@ 1.1. Motivation
The Adler–Kostant–Symes (AKS) Theorem is a fundamental Theorem in the theory of Hamiltonian systems. It allows to associate to a splitting of a finite-dimensional Lie algebra a new Lie bracket leading to isospectral evolutions, i.e. with spectral functions in involutions. Although the AKS theorem and similar involutivity Theorems have been extensively applied in the infinite-dimensional setting (see [Adl78, Sym80a, Sym80b, Rat80, BMR13, CP24] and the references therein), a non-formal presentation in the Banach setting seems to be lacking. Recent applications to the theory of Lie Group Thermodynamic [Bar25] motivate us to think about the foundations of this theory in the Banach case.
@@ 1.2. Structure of the paper
The first section contains a summary of the theory of Banach Poisson–Lie groups developed in [Tum20]. It is as self-contained as possible and can be used as a first introduction to the subject. In particular, the equivalence between Manin triples and Banach Lie–Poisson spaces which are Banach Lie bialgebras is presented. This equivalence is at the heart of the necessity to extend the notion of Banach Poisson manifolds to the one presented in section 2.7. In sections 3, 4 and 5 we recall different approaches that lead to the definition of an auxiliary Lie bracket on a Banach Lie algebra using an operator satisfying some equations, like the modified Classical Yang-Baxter equation, the Baxter equation or the vanishing of the Nijenhuis torsion. This auxiliary bracket can lead to the existence of a new structure of Banach Lie–Poisson space on any space in duality with the original Banach Lie algebra. Section 3.3 contains the involutivity theorems that we prove in the Banach context. In section 6, the equations of motion on coadjoint orbits are transported to adjoint orbits using an $\operatorname{\textnormal{Ad}}_{G}$-invariant pairing, leading to equations in Lax form. The resolution of these equations using the solution of the factorization problem is presented and applied to the Iwasawa decomposition. In section 7.3, we present how the theory allows to recover the equations of the semi-infinite Toda lattice in Flaschka coordinates.
@@ 2.1. Notions of Poisson manifolds in the Banach setting
To the best of our knowledge it is not even known if Poisson brackets need to be localizable, i.e. depend only on the germ of functions at a particular point, see [BGT18]. In the finite dimensional case this fact follows from Leibniz property and the existence of bump functions (i.e. non-zero functions with compact support). However on Banach manifolds (or even on Banach spaces) there may be no bump functions, see [BF66] or discussion in [CP12].
@@ 2.2. Banach Lie–Poisson spaces
In general, a closed subspace of a Banach space admitting a predual might not admit a predual. For instance the subspace of compact operators on a Hilbert space is a closed subspace of the Banach space of bounded operators which does not admit a predual, whereas the Banach space of bounded operator has the space of trace class operators as predual. Even if the predual does exist, it might not be unique and it is not guaranteed that it will be preserved by coadjoint action. Thus if $\mathfrak{g}$$_{*}$ is a Banach Lie–Poisson space predual to $\mathfrak{g}$}, and ${\mathfrak{g}}_{+}\subset{\mathfrak{g}}$ is a closed Lie subalgebra, there might not be a Banach Lie–Poisson space predual to $\mathfrak{g}$$_{+}$.
@@ 2.3. Generalized Banach Poisson manifolds
Then $\{\cdot,\cdot\}:\mathcal{A}\times\mathcal{A}\rightarrow\mathcal{A}$ satisfies conditions 2.1 and is called a generalized Poisson bracket on $M$.
@@ 2.5. Banach Lie–Poisson spaces for an arbitrary duality pairing
Suppose that a Banach Lie algebra $\mathfrak{g}$} of a Banach Lie group $G$ admits a non-degenerate continuous bilinear map $\langle\cdot,\cdot\rangle:{\mathfrak{g}}\times{\mathfrak{g}}\rightarrow{\mathfrak{g}}$, invariant by the adjoint action of $\mathfrak{g}$}, and denote by $\iota:{\mathfrak{g}}\rightarrow{\mathfrak{g}}^{*}$ the injection which maps $X\in{\mathfrak{g}}$ to $\langle X,\cdot\rangle\in{\mathfrak{g}}^{*}$. Then $\mathfrak{g}$} is a Banach Lie–Poisson space with respect to itself. The Hamiltonian vector field associated to a smooth function $H$ in $\mathcal{A}$ is given by
@@ 3.2. $R$-matrices associated with the sum of Banach Lie subalgebras
We shall present now a widely used method of obtaining examples of classical $R$-matrices, namely when the Lie algebra $\mathfrak{g}$ admits a Banach decomposition into the direct sum of two closed Lie Banach subalgebras: $\mathfrak{g}=\mathfrak{g}_{+}\oplus\mathfrak{g}_{-}$. This situation can be traced back under different names in the literature: under the name “twilled extension” or “twilled Lie algebra” [KS89, KSM88], or “algèbre de Lie bicroisée” [Ami88], under the name bicrossproduct Lie algebra [Maj88], or under the name double Lie algebra in [LW90], which differs from the more general definition of double Lie algebra given in Definition 3.1.
@@ 3.3.1. Involutivity Theorem (simplified version of AKS theorem with $\varepsilon=0$)
Recall that, for a Banach Lie–Poisson space $\mathfrak{b}$} with respect to a Banach Lie algebra $\mathfrak{g}$}, $\mathcal{A}$ denotes the unital subalgebra of $\mathscr{C}^{\infty}({\mathfrak{b}})$ consisting of all functions with differentials in $\mathfrak{g}$}, see (2.11). Let us first present a simplified version of the AKS theorem (with argument shift $\varepsilon=0$), but suitable to our Banach setting of generalized Poisson structures. We refer the reader to [STS83, Theorem 1] or [AvMV04, Theorem 4.36] for the original versions of the following theorem in finite dimensional setting. Note that in points (1) and (2) of Theorem 3.9, we don’t assume that the $R$-matrix comes from the decomposition of the Banach Lie algebra $\mathfrak{g}$}.
@@
for all $x\in{\mathfrak{g}}_{\mp}$ and $b\in{\mathfrak{g}}_{\pm}^{0}\cap{\mathfrak{b}}$, and $\operatorname{\textnormal{ad}}^{*}:{\mathfrak{g}}_{\mp}\times{\mathfrak{g}}_{\pm}^{0}\cap{\mathfrak{b}}\rightarrow{\mathfrak{g}}_{\pm}^{0}\cap{\mathfrak{b}}$ is continuous. The fact that the coadjoint action of $\mathfrak{g}$$_{\mp}$ preserves ${\mathfrak{g}}_{\pm}^{0}\cap{\mathfrak{b}}$ follows from the fact that the coadjoint action of $\mathfrak{g}$$_{\mp}$ preserves ${\mathfrak{g}}_{\mp}^{*}\simeq{\mathfrak{g}}_{\pm}^{0}$, and also $\mathfrak{b}$}, since $\mathfrak{b}$} is a Banach–Lie Poisson space with respect to $\mathfrak{g}$}. The continuity of the coadjoint actions follow from the continuity of the coadjoint action of $\mathfrak{g}$} on $\mathfrak{b}$} and of the projections. ∎
@@ 4.2. Rota-Baxter Lie groups
    The Lie group $SL(n,{\mathbb{C}})$ of complex $n\times n$ matrices with determinant $1$ factorizes as
@@ 5.1. Linear Nijenhuis Operators and associated Lie brackets
Nijenhuis Operators in the Banach setting where introduced in [GLT24a] and used in [GLT24b] in relation to $C^{*}$-algebras. In this section, we recall the Lie bracket associated to a Nijenhuis operator and the relation to Rota-Baxter algebras. We follow the presentation given in [KSM90] for the finite-dimensional case. This section will be applied to the semi-infinite Toda lattice in section 7.3.
@@ 6. Lax equations associated with Banach–Poisson Lie groups
$\frac{d}{dt}_{|t=0}\operatorname{\textnormal{Ad}}^{*}_{g(t)}\mu=\operatorname{\textnormal{ad}}^{*}_{M}\mu$
@@
By equation (6.2),
$\operatorname{\textnormal{ad}}^{*}_{M}\mu(Y)=\langle L,[M,Y]\rangle=-\langle[M,L],Y\rangle.$
Consequently
@@ 7.3. Semi-infinite Toda lattice
$\dot{\mathbf{q}}_{n}$ $=\frac{1}{2}\left({\mathbf{p}}_{n-1}\frac{\partial H}{\partial{\mathbf{p}}_{n-1}}-{\mathbf{p}}_{n}\frac{\partial H}{\partial{\mathbf{p}}_{n}}\right)$
@@
$\dot{\mu}=-(\operatorname{\textnormal{ad}}^{*}_{N})_{x_{{\mathbf{a}}{\mathbf{b}}}}\mu$
@@
$\dot{\mu}=-\frac{1}{2}\big(\operatorname{\textnormal{ad}}^{*}_{Nx_{{\mathbf{a}}{\mathbf{b}}}}+[N^{*},\operatorname{\textnormal{ad}}^{*}_{x_{{\mathbf{a}}{\mathbf{b}}}}]\big)\mu=-\frac{1}{2}p_{L^{q}(\mathcal{H})_{-}}\big([Nx_{{\mathbf{a}}{\mathbf{b}}},\mu]+N^{*}[x_{{\mathbf{a}}{\mathbf{b}}},\mu]-[x_{{\mathbf{a}}{\mathbf{b}}},N^{*}\mu]\big)$
@@
$2\dot{\mathbf{q}}$ $=-\delta^{-}({\mathbf{b}}\cdot{\mathbf{p}})$
$2\dot{\mathbf{p}}$ $={\mathbf{p}}\cdot\delta^{+}{\mathbf{a}}$
@@
$2\dot{\mathbf{q}}_{n}$ $={\mathbf{b}}_{n}{\mathbf{p}}_{n}-{\mathbf{b}}_{n-1}{\mathbf{p}}_{n-1}={\mathbf{p}}_{n}\frac{\partial H}{\partial{\mathbf{p}}_{n}}-{\mathbf{p}}_{n-1}\frac{\partial H}{\partial{\mathbf{p}}_{n-1}}$
@@
$\dot{\mathbf{q}}_{n}$ $=2({\mathbf{p}}_{n}^{2}-{\mathbf{p}}_{n-1}^{2})$

### v2

@@ 1.1. Motivation
The Adler–Kostant–Symes (AKS) theorem is a fundamental theorem in the theory of Hamiltonian systems. It allows to associate to a splitting of a finite-dimensional Lie algebra a new Lie bracket leading to isospectral evolutions, i.e. with spectral functions in involutions. Although the AKS theorem and similar involutivity theorems have been extensively applied both in the finite and infinite-dimensional setting (see [Adl78, Sym80a, Sym80b, Rat80, BMR13, CP24] and the references therein), a non-formal presentation in the Banach setting seems to be lacking. Recent applications to the theory of Lie group thermodynamic [Bar25] motivate us to think about the foundations of this theory in the Banach case.
@@ 1.2. Structure of the paper
The first section contains a summary of the theory of Banach Poisson–Lie groups developed in [Tum20]. It is as self-contained as possible and can be used as a first introduction to the subject. In particular, the equivalence between Manin triples and Banach Lie–Poisson spaces which are Banach Lie bialgebras is presented. This equivalence is at the heart of the necessity to extend the notion of Banach Poisson manifolds to the one presented in section 2.7. In sections 3, 4 and 5 we recall different approaches that lead to the definition of an auxiliary Lie bracket on a Banach Lie algebra using an operator satisfying some equations, like the modified classical Yang–Baxter equation, the Baxter equation or the vanishing of the Nijenhuis torsion. This auxiliary bracket can lead to the existence of a new structure of Banach Lie–Poisson space on any space in duality with the original Banach Lie algebra. Section 3.3 contains the involutivity theorems that we prove in the Banach context. In section 6, the equations of motion on coadjoint orbits are transported to adjoint orbits using an $\operatorname{\textnormal{Ad}}_{G}$-invariant pairing, leading to equations in Lax form. The resolution of these equations using the solution of the factorization problem is presented and applied to the Iwasawa decomposition. In section 7.3, we present how the theory allows to recover the equations of the semi-infinite Toda lattice in Flaschka coordinates.
@@ 2.1. Notions of Poisson manifolds in the Banach setting
To the best of our knowledge it is not even known if Poisson brackets need to be localizable, i.e. depend only on the germ of functions at a particular point, see [BGT18]. In the finite dimensional case this fact follows from Leibniz property and the existence of bump functions (i.e. non-zero functions with compact support). However on Banach manifolds (or even on Banach spaces) there may be no bump functions, see [Kur54, BF66] or discussion in [CP12].
@@ 2.2. Banach Lie–Poisson spaces
In general, a closed subspace of a Banach space admitting a predual might not admit a predual. For instance the subspace of compact operators on a Hilbert space is a closed subspace of the Banach space of bounded operators which does not admit a predual, whereas the Banach space of bounded operator has the space of trace class operators as predual. Even if the predual does exist, it might not be unique and it is not guaranteed that it will be preserved by coadjoint action. Thus if $\mathfrak{g}$$_{*}$ is a Banach Lie–Poisson space predual to $\mathfrak{g}$}, and ${\mathfrak{g}}_{+}\subset{\mathfrak{g}}$ is a closed Lie subalgebra, there might not be a Banach Lie–Poisson space predual to $\mathfrak{g}$$_{+}$. See also the discussion in the context of precotangent bundles in [Gol25].
@@ 2.3. Generalized Banach Poisson manifolds
Then $\{\cdot,\cdot\}:\mathcal{A}\times\mathcal{A}\rightarrow\mathcal{A}$ satisfies conditions (i)(iii)(i)-(iii) from Definition 2.1 and is called a generalized Poisson bracket on $M$.
@@ 2.5. Banach Lie–Poisson spaces for an arbitrary duality pairing
Suppose that a Banach Lie algebra $\mathfrak{g}$} of a Banach Lie group $G$ admits a non-degenerate continuous bilinear map $\langle\cdot,\cdot\rangle:{\mathfrak{g}}\times{\mathfrak{g}}\rightarrow{\mathfrak{g}}$, invariant by the adjoint action of $\mathfrak{g}$}, and denote by $\iota:{\mathfrak{g}}\rightarrow{\mathfrak{g}}^{*}$ the injection which maps $X\in{\mathfrak{g}}$ to $\langle X,\cdot\rangle\in{\mathfrak{g}}^{*}$. Then $\mathfrak{g}$} is a Banach Lie–Poisson space with respect to itself.
The Hamiltonian vector field associated to a smooth function $H$ in $\mathcal{A}$
is given by
@@ 3.2. $R$-matrices associated with the sum of Banach Lie subalgebras
We shall present now a widely used method of obtaining examples of classical $R$-matrices, namely when the Lie algebra $\mathfrak{g}$ admits a Banach decomposition into the direct sum of two closed Lie Banach subalgebras: $\mathfrak{g}=\mathfrak{g}_{+}\oplus\mathfrak{g}_{-}$. This situation can be traced back under different names in the literature: under the name “twilled extension” or “twilled Lie algebra” [KS89, KSM88], or “algèbre de Lie bicroisée” [Ami88], under the name “bicrossproduct Lie algebra” [Maj88], or under the name “double Lie algebra” in [LW90], which differs from the more general definition of double Lie algebra given in Definition 3.1.
@@ 3.3.1. Involutivity theorem (simplified version of AKS theorem with $\varepsilon=0$)
Recall that, for a Banach Lie–Poisson space $\mathfrak{b}$} with respect to a Banach Lie algebra $\mathfrak{g}$}, $\mathcal{A}$ denotes the unital subalgebra of $\mathscr{C}^{\infty}({\mathfrak{b}})$ consisting of all functions with differentials in $\mathfrak{g}$}, see (2.11).
Let us first present a simplified version of the AKS theorem (with argument shift $\varepsilon=0$), but suitable to our Banach setting of generalized Poisson structures. We refer the reader to [STS83, Theorem 1] or [AvMV04, Theorem 4.36] for the original versions of the following theorem in finite dimensional setting. Note that in points (1) and (2) of Theorem 3.9, we don’t assume that the $R$-matrix comes from the decomposition of the Banach Lie algebra $\mathfrak{g}$}.
@@
for all $x\in{\mathfrak{g}}_{\mp}$ and $b\in{\mathfrak{g}}_{\pm}^{0}\cap{\mathfrak{b}}$, and $\operatorname{\textnormal{ad}}^{*}:{\mathfrak{g}}_{\mp}\times{\mathfrak{g}}_{\pm}^{0}\cap{\mathfrak{b}}\rightarrow{\mathfrak{g}}_{\pm}^{0}\cap{\mathfrak{b}}$ is continuous. The fact that the coadjoint action of $\mathfrak{g}$$_{\mp}$ preserves ${\mathfrak{g}}_{\pm}^{0}\cap{\mathfrak{b}}$ follows from the fact that the coadjoint action of $\mathfrak{g}$$_{\mp}$ preserves ${\mathfrak{g}}_{\mp}^{*}\simeq{\mathfrak{g}}_{\pm}^{0}$, and also $\mathfrak{b}$}, since $\mathfrak{b}$} is a Banach–Lie Poisson space with respect to $\mathfrak{g}$}. The continuity of the coadjoint actions follows from the continuity of the coadjoint action of $\mathfrak{g}$} on $\mathfrak{b}$} and of the projections. ∎
@@ 4.2. Rota-Baxter Lie groups
    Lie group $SL(n,{\mathbb{C}})$ of complex $n\times n$ matrices with determinant $1$ factorizes as
@@ 5.1. Linear Nijenhuis operators and associated Lie brackets
Nijenhuis operators were applied in the theory of integrable, see e.g. [KSM90, BKM22, KS19]. In the Banach setting they were studied in [GLT24a] and used in [GLT24b, GG25]. In this section, we recall the Lie bracket associated to a Nijenhuis operator and the relation to Rota-Baxter algebras. We follow the presentation given in [KSM90] for the finite-dimensional case. The results of this section will be applied to the semi-infinite Toda lattice in section 7.3.
@@ 6. Lax equations associated with Banach–Poisson Lie groups
$\frac{d}{dt}_{|t=0}\operatorname{\textnormal{Ad}}^{*}_{g(t)}\mu=\operatorname{\textnormal{ad}}^{*}_{M}\mu,$
@@
By equation (6.2), we get
$\operatorname{\textnormal{ad}}^{*}_{M}\mu(Y)=\langle L,[M,Y]\rangle=-\langle[M,L],Y\rangle,$
which can be also written as
@@ 7.3. Semi-infinite Toda lattice
$\dot{\mathbf{q}}_{n}$ $=\frac{1}{2}\left({\mathbf{p}}_{n-1}\frac{\partial H}{\partial{\mathbf{p}}_{n-1}}-{\mathbf{p}}_{n}\frac{\partial H}{\partial{\mathbf{p}}_{n}}\right),$
@@
$\dot{\mu}=-(\operatorname{\textnormal{ad}}^{*}_{N})_{x_{{\mathbf{a}}{\mathbf{b}}}}\mu.$
@@
$\dot{\mu}=-\frac{1}{2}\big(\operatorname{\textnormal{ad}}^{*}_{Nx_{{\mathbf{a}}{\mathbf{b}}}}+[N^{*},\operatorname{\textnormal{ad}}^{*}_{x_{{\mathbf{a}}{\mathbf{b}}}}]\big)\mu=-\frac{1}{2}p_{L^{q}(\mathcal{H})_{-}}\big([Nx_{{\mathbf{a}}{\mathbf{b}}},\mu]+N^{*}[x_{{\mathbf{a}}{\mathbf{b}}},\mu]-[x_{{\mathbf{a}}{\mathbf{b}}},N^{*}\mu]\big).$
@@
$2\dot{\mathbf{q}}$ $=-\delta^{-}({\mathbf{b}}\cdot{\mathbf{p}}),$
$2\dot{\mathbf{p}}$ $={\mathbf{p}}\cdot\delta^{+}{\mathbf{a}}.$
@@
$2\dot{\mathbf{q}}_{n}$ $={\mathbf{b}}_{n}{\mathbf{p}}_{n}-{\mathbf{b}}_{n-1}{\mathbf{p}}_{n-1}={\mathbf{p}}_{n}\frac{\partial H}{\partial{\mathbf{p}}_{n}}-{\mathbf{p}}_{n-1}\frac{\partial H}{\partial{\mathbf{p}}_{n-1}},$
@@
$\dot{\mathbf{q}}_{n}$ $=2({\mathbf{p}}_{n}^{2}-{\mathbf{p}}_{n-1}^{2}),$